# Filter Like You Test: Data-Driven Data Filtering for CLIP Pretraining

### v2

@@ C.1 Experiment configurations
Upstream batch size ${B}$ 4096 4096 4096 4096
@@ C.2 Computational cost
In terms of FLOPs, M-FLYT requires slightly more than 2×2$\times$2 × the FLOPs of regular CLIP model training per example, while E-FLYT requires slightly more than 3×3$\times$3 ×. Since we only train our models for 20M examples (compared to DataComp’s 128M), M-FLYT uses approximately 3×3$\times$3 × fewer FLOPs than training a DataComp medium scale model, and E-FLYT uses approximately 2×2$\times$2 × fewer.
With our implementation, M-FLYT takes 40 A100 hours to train, and E-FLYT takes 45 A100 hours, roughly equivalent to running a DataComp medium scale experiment. For SCS with a batch size ${G}=100K$, it takes about 100 minutes to sample 128M examples using a naive Python implementation on a single CPU. For larger scales (DataComp large or Xlarge), the filtering computation becomes negligible relative to the overall training cost.
@@ C.5 M-FLYT baseline - ImageNet-weighted standardized sum
We then normalized the ImageNet accuracies $\mathrm{IN}_{1:K}$ between zero and one, and add a scalar to achieve our desired ratio ${r}$:
@@
where $\mathrm{min}(\mathrm{IN}_{1:K})$ is the lowest ImageNet score, and $\mathrm{max}(\mathrm{IN}_{1:K})$ is the highest. This ensures that the ratio between the maximum and minimum weights equals ${r}$:
@@ C.10 SCS ablations
With SCS, we evaluate different repetition penalties and find that performance varies smoothly with ${\alpha}$, reaching optimal results around ${\alpha}=0.15$ (Table 13). For sampling efficiency, we introduce the batch size ${G}$, setting it to $100K$. This value is large enough for efficient medium scale sampling yet sufficiently small based on our empirical observation that while each sample had 1280 ($128M/100K$) opportunities to be selected, our most frequently sampled example appears only 56 times. To more directly validate our choice of ${G}$, Table 13 shows that using values of ${G}$ of different order of magnitude has little effect on performance.
@@
Repetition penalty ${\alpha}{}$ Batch size ${G}{}$ ImageNet Average
@@ C.11 DataComp small scale results
For the small scale DataComp benchmark, we reuse the mixing model from the medium scale without retraining E-FLYT or M-FLYT. The only changes were re-tuning the repetition penalty ${\alpha}$ for the smaller dataset and reducing the sampling batch size ${G}$ by a factor of 10 to $10K$, aligning with the dataset size reduction. Figure 4 shows that the smaller scale performs best with a higher repetition penalty, around ${\alpha}=0.5$, compared to the medium scale which performs better with values between 0.15 to 0.25. A higher repetition penalty reduces the number of repetitions per example, which becomes more important for smaller datasets.

### v3

@@ C.1 Experiment configurations
Upstream batch size B{B} 4096 4096 4096 4096
@@ C.2 Computational cost
In terms of FLOPs, M-FLYT requires slightly more than $2\times$ the FLOPs of regular CLIP model training per example, while E-FLYT requires slightly more than $3\times$. Since we only train our models for 20M examples (compared to DataComp’s 128M), M-FLYT uses approximately $3\times$ fewer FLOPs than training a DataComp medium scale model, and E-FLYT uses approximately $2\times$ fewer.
With our implementation, M-FLYT takes 40 A100 hours to train, and E-FLYT takes 45 A100 hours, roughly equivalent to running a DataComp medium scale experiment. For SCS with a batch size G=100K{G}=100K, it takes about 100 minutes to sample 128M examples using a naive Python implementation on a single CPU. For larger scales (DataComp large or Xlarge), the filtering computation becomes negligible relative to the overall training cost.
@@ C.5 M-FLYT baseline - ImageNet-weighted standardized sum
We then normalized the ImageNet accuracies $\mathrm{IN}_{1:K}$ between zero and one, and add a scalar to achieve our desired ratio r{r}:
@@
where $\mathrm{min}(\mathrm{IN}_{1:K})$ is the lowest ImageNet score, and $\mathrm{max}(\mathrm{IN}_{1:K})$ is the highest. This ensures that the ratio between the maximum and minimum weights equals r{r}:
@@ C.10 SCS ablations
With SCS, we evaluate different repetition penalties and find that performance varies smoothly with $\alpha$}, reaching optimal results around ${\alpha}=0.15$ (Table 14). For sampling efficiency, we introduce the batch size G{G}, setting it to $100K$. This value is large enough for efficient medium scale sampling yet sufficiently small based on our empirical observation that while each sample had 1280 ($128M/100K$) opportunities to be selected, our most frequently sampled example appears only 56 times. To more directly validate our choice of G{G}, Table 14 shows that using values of G{G} of different order of magnitude has little effect on performance.
@@
Repetition penalty $\alpha$}{} Batch size G{G}{} ImageNet Average
@@ C.11 DataComp small scale results
For the small scale DataComp benchmark, we reuse the mixing model from the medium scale without retraining E-FLYT or M-FLYT. The only changes were re-tuning the repetition penalty $\alpha$} for the smaller dataset and reducing the sampling batch size G{G} by a factor of 10 to $10K$, aligning with the dataset size reduction. Figure 4 shows that the smaller scale performs best with a higher repetition penalty, around ${\alpha}=0.5$, compared to the medium scale which performs better with values between 0.15 to 0.25. A higher repetition penalty reduces the number of repetitions per example, which becomes more important for smaller datasets.
@@ Appendix D Reproduction comparison
Appendix E Qualitative analysis
Figure 5: M-FLYT top scoring examples.
Figure 6: CLIP ViT-B/32 top scoring examples.
Figure 7: DFN-FT top scoring examples.
Figures 5, 6 and 7 show the top-scoring examples for M-FLYT, standard CLIP score, and DFN-FT. We observe that M-FLYT prioritizes examples containing multiple ImageNet classes. For instance, the highest-scoring example features the caption ”Honey the Golden Retriever stands in a tent,” where both ”Golden Retriever” and ”tent” correspond to ImageNet classes. In contrast, the CLIP score filtering method using OpenAI’s ViT-B/32 model favors images with visible text that matches the caption text.